# Zero-Shot Large Language Models for Long Clinical Text Summarization with Temporal Reasoning

### v2

@@ Abstract
Recent advances in large language models (LLMs) have shown potential in clinical text summarization, but their ability to handle long patient trajectories with multi-modal data spread across time remains underexplored. This study systematically evaluates several state-of-the-art open-source LLMs and their Retrieval Augmented Generation (RAG) variants on long-context clinical summarization. We examine their ability to synthesize structured and unstructured Electronic Health Records (EHR) data while reasoning over temporal coherence, by re-engineering existing tasks, including discharge summarization and diagnosis prediction from two publicly available EHR datasets. Our results indicate that long context window improves input integration, but do not consistently enhance clinical reasoning, and LLMs are still struggling with temporal progression and rare disease prediction. While RAG show improvements in hallucination in some cases, it does not fully address these limitations. Our work fills the gap in long clinical text summarization, establishing a foundation for evaluating LLMs with multi-modal data and temporal reasoning.
Zero-Shot Large Language Models for Long Clinical Text Summarization with Temporal Reasoning
@@ 1 Introduction
Electronic Health Records (EHRs) encapsulate a wide range of multi-modal data, such as vital signs, laboratory results, radiology findings, and free text clinical notes Mohsen et al. (2022); Li et al. (2022); Belden et al. (2017). Organized across various timestamps, they reflect the dynamic nature of patient care. In clinical settings, particularly for older patients in intensive care units (ICUs) with multiple encounters, chronic conditions, and complex treatment plans, EHRs become especially lengthy and intricate. Clinicians reviewing these long patient histories face significant cognitive burden, often leading to inaccurate decisions such as diagnostic errors Dymek et al. (2021); Singh et al. (2017). Automated summarization can help improve care continuity and decision-making Dymek et al. (2021); Adams et al. (2021); Gao et al. (2023a); Laxmisan et al. (2012); Pivovarov and Elhadad (2015); Liang et al. (2019), but the volume and complexity of EHRs pose challenges, requiring models that handle diverse and evolving clinical data. As in Figure 1, summarizing longitudinal patient trajectories involves extracting temporally ordered clinical events from multi-modal sources. In critical care, robust temporal reasoning is essential to preserve causal relationships among medical events.
@@
Our work evaluates large language models (LLMs) for summarizing complex, longitudinal EHRs, capturing full patient trajectories rather than isolated clinical snapshots. While LLMs have shown promise in clinical NLP tasks Silcox et al. (2024); Wachter and Brynjolfsson (2024); Garcia et al. (2024); Adams et al. (2024); Gao et al. (2022), existing evaluations primarily focus on short-context settings and task-specific fine-tuning. Clinical summarization benchmarks, such as discharge summarization (“Discharge Me”Xu et al. (2024a)) and diagnosis generation from progress notes (“ProbSum”Gao et al. (2023a)), evaluate LLMs on isolated segments of patient records rather than full hospital trajectories.
While techniques like Retrieval-Augmented Generation (RAG) Xu et al. ; Gao et al. (2023b); Lewis et al. (2020) have been explored for diagnosis prediction and discharge summarization Myers et al. (2024); Lyu et al. (2024), their effectiveness in processing long clinical narratives with rich temporal dependencies remains underexplored.
Two key gaps motivate our work: 1) As foundational LLMs with long-context capabilities (over 4k tokens) are adopted into clinical workflows, it is crucial to evaluate their performance without task-specific training to understand their inherent strengths and limitations in processing complex medical information. (2) Existing task settings focus on single time-point records, failing to reflect real-world longitudinal patient care. Our study addresses these gaps by systematically assessing LLMs on full patient trajectories, providing insights into their ability to manage temporally evolving clinical narratives and their applicability in real-world settings.
@@
In this paper, we focus on two public EHR dataset: Medical Information Mart for Intensive Care (MIMIC-III) Johnson et al. (2020) and EHRShot Wornow et al. (2023). We evaluate five state-of-the-art LLMs and their RAG variants: Mistral-7B-Instruct-v0.1  Jiang et al. (2023), Llama3-8B-Instruct  AI@Meta (2024), Qwen2.5-7B  Yang et al. (2024), DeepSeek-R1-Distill-Qwen-32B  DeepSeek-AI et al. (2025) and Llama2-13B-chat-hf  Touvron et al. (2023). Our work contributes to the LLM for clinical text summarization through the following aspects:
@@
  • We re-engineer existing tasks, including discharge summarization, progress note summarization, and diagnosis classification, into new settings for long-context summarization with temporal reasoning, explicitly designed to align with real-world clinical workflows.
  • The tasks allow us to systematically investigate the effects of different modalities and timestamps by incorporating both unstructured clinical notes and structured tabular data.
  • We compare different methods for handling long documents, including directly processing EHR inputs with LLMs and using RAG approaches.
Given the absence of a benchmark for long clinical text summarization that integrates multi-modal data, temporal reasoning, and real-world clinical workflows, our work establishes a foundation for evaluating LLMs in this setting. We provide insights into their strengths and limitations while guiding future research toward more comprehensive, clinically-aligned, and temporally-aware summarization models.
@@ 2 Related Work
Directly processing the long document input can lead to LLMs’ “lost-in-the-middle" problem Liu et al. (2024). As a result, RAG has been the mainstream when handling long clinical documents, evidenced by their superior performances in diagnosis prediction, discharge summarization and information extraction Myers et al. (2024); Lyu et al. (2024); Lopez et al. (2025). Due to the task setup, these works have only investigated single modality EHRs with long input length.
Our work addresses this problem by extending RAG-based approaches to multi-modal EHR summarization, incorporating both structured (tabular) and unstructured (clinical notes) data to enhance the completeness and accuracy of patient trajectory summaries. We evaluate how well LLMs and their RAG set up can capture and preserve temporal and causal relationships across the full course of a patient’s hospital stay, bridging the gap between isolated document processing and comprehensive longitudinal understanding.
@@ 3.1 MIMIC-III
MIMIC-III contains data from over 40,000 patients admitted to the ICU at Beth Israel Deaconess Medical Center between 2001 and 2012. Out of all patients, we decided to focus on older patients (65 years and older) who have been admitted to the ICU, accounting for 43.09% of dataset’s total patient population. After consulting with clinicians, we selected patients with a hospital stay exceeding 72 hours from admission, ensuring sufficient accumulated information for summarization while allowing medical uncertainty to stabilize. This specially selected cohort sets our work apart from prior MIMIC-III studies by focusing on older ICU patients who benefit most from accurate LLM-driven summarization, addressing complex cases where clear summaries are crucial for decision-making. By further restricting our selection to hospital stays exceeding 72 hours, we enable a rigorous evaluation of LLMs’ ability to summarize long documents and reason over progressive clinical events, ultimately supporting better care for a population at high risk of adverse outcomes.
MIMIC-III includes physician and nursing notes, ECG reports, and other clinical documents, along with a large amount of structured (tabular) data. We focus on chart events, lab events, input events, and medications, as they are most relevant to patient care. Chart events capture routine vital signs and other bedside measurements, including ventilator settings and mental status. Lab events consist of laboratory test results, such as white blood cell count and cholesterol levels. Input events track administered fluids, including oral/tube feedings and IV infusions, while medications record all drugs given to the patient.
3.1.1 Discharge Summarization
@@
3.1.2 Assessment and Plan (A&P) Generation
@@
The setup of the task is illustrated in Figure 2, which explains the three input methods: the no prior context (Baseline) method uses only the current day’s structured EHR data, the Single-Day Context method adds the previous day’s progress note for limited historical context, and the Multi-Day Context method incorporates all prior progress notes for a more comprehensive patient history. For day $i$, the input includes progress notes up to $i-1$ (depending on the method), but the note for $i$ is excluded and compared to the generated version. This setup reflects the real clinical workflow, where physicians reference prior notes when writing new ones.
@@
Figure 3 shows the input token lengths for both tasks across different settings. Discharge summarization has an average of 500 tokens from the 24 hour and  1940 tokens from the 48 hour window. On Assessment and Plan (A&P) generation, the average token length increases significantly, due to the need to incorporate prior progress notes and structured data over multiple days. The No Context (Baseline) method starts with a relatively high token count on Day 1 but remains lower overall, averaging around 3,125 tokens per day. The Single-Day Context method, which incorporates the previous day’s progress note, leads to a moderate increase in token length, averaging 5,375 tokens. In contrast, the Multi-Day Context method, which accumulates all prior notes, shows the most significant token growth, reaching nearly 10,000 tokens on Day 4 and averaging 6,875 tokens.
@@ 3.2 Diagnosis Prediction from EHRShot
The EHRShot dataset  Wornow et al. (2023) is a longitudinal dataset containing only structured EHR data from 6,739 patients at Stanford Medicine, with over 41 million clinical events. Its structured nature makes it distinct from MIMIC, allowing for a more controlled evaluation of models’ ability to process structured temporal information. Additionally, EHRShot includes a broad range of hospital encounters, making it a more comprehensive resource for modeling long-term patient trajectories.
Although our primary focus is on summarizing longitudinal and temporal EHR data, we chose to evaluate models on the six binary diagnosis prediction tasks available in EHRShot. These tasks involve predicting whether a patient will receive a given diagnosis within one year after discharge. While the final prediction is binary, the decision-making process requires synthesizing longitudinal patient history across multiple clinical events, a critical capability for summarization models designed to handle temporal reasoning in EHR data. The selected diagnoses and their prevalence in the dataset are: Hypertension (28.64%), Hyperlipidemia (31.85%), Pancreatic Cancer (20.49%), Celiac (3.46%), Lupus (4.44%), and Acute MI (Myocardial Infarction) (37.04%). The input token lengths vary widely, with a mean of 1,989 tokens ($\sigma$: 1,270) and a median of 1,851. For distinct measurements per patient, the mean is 74 ($\sigma$:48) and a median of 66. More statistics are presented in Appendix A.2.
@@ 3.3 Representing multi-modal, longitudinal patient data
To support the LLMs’ understanding of the structured data, we transform it into natural language, following previous table-to-text manners Gao et al. (2024); Yu et al. (2023). As a first step, the codes for the various measurements in the data are replaced by their natural language description and embedded in a sentence in the form of [Measurement] is [Value][Unit of Measurement]. Input events and medications use the phrase ‘[Measurement] is administered’. Then, after the temporal ordering, the absolute datetime values are replaced by relative values with regard to the preceding timestamp. Measurements taken at the same time are grouped together under one timestamp. An example of a patient chronology consisting of structured data converted to narrative format with a radiology note is given in Table 1.
It is well known that clinical information is frequently recorded and copy-pasted multiple times, a phenomenon commonly referred to as "note bloat," which leads to redundancy and makes it more challenging to extract relevant insights  Shoolin et al. (2013); Liu et al. (2022). This issue is further exacerbated by the large volume of data, which poses challenges when working with LLMs due to their limited context windows. To address this problem and reduce redundancy, we apply preprocessing steps to filter out duplicate records. If a measurement appears in both lab events and chart events, we retain only one instance to prevent duplication. Additionally, when multiple identical records exist within a short time frame, we always retain the latest entry. This approach applies to both structured data and clinical notes, ensuring that copy-pasted or redundant information do not inflate the dataset unnecessarily. On EHRShot data, we remove redundant words from the constructed patient chronologies, and organize data into six-hour chunks based on their timestamp. Similarly to our MIMIC patient cohort, we also focus on patients aged 65 and above, as well as filter out patients with less than 30 measurements. Other details are included in Appendix A.1.
@@ 4 Methods
We test five LLMs with various context length: Llama2-13B-chat-hf Touvron et al. (2023) and Mistral-7B-Instruct-v0.1 Jiang et al. (2023), both with a context window of 4K. Additionally, we evaluate Llama3-8B-Instruct AI@Meta (2024), which supports an 8K context window. For LLMs with larger context windows, we test Qwen2.5-7B Yang et al. (2024) and DeepSeek-R1-Distill-Qwen-32B DeepSeek-AI et al. (2025), both capable of handling 128K token input.
@@ 4.1 Approaches
We compare a baseline direct generation approach and a structured RAG approach. The key difference between these approaches lies in how they handle long temporal contexts. Direct generation processes the entire input at once but may suffer from the lost-in-the-middle problem, where relevant information buried in long documents is overlooked. In contrast, RAG segments temporal information into retrievable chunks, which helps manage long sequences but may disrupt temporal dependencies between events.
Direct Generation
@@
Experiment settings for Direct Generation and RAG
For all LLMs, we run their 8-bit quantized version. We set the output token length as 1,000, but almost all tasks output is significantly shorter than this limit. For RAG setup, we used Langchain Chase (2022) with Faiss Douze et al. (2024) for semantic retrieval in vector database. We perform hyperparameter tuning with chunk size between [250, 750], top k between [10, 50], chunk overlap between [50,200]. The detailed results of hyperparameter searching are in Appendix A.3. All experiments are run on 2 H100 94GB GPUs.

### v3

# Large Language Models with Temporal Reasoning for Longitudinal Clinical Summarization and Prediction

@@ Abstract
Recent advances in large language models (LLMs) have shown potential in clinical text summarization, but their ability to handle long patient trajectories with multi-modal data spread across time remains underexplored. This study systematically evaluates several state-of-the-art open-source LLMs, their Retrieval Augmented Generation (RAG) variants and chain-of-thought (CoT) prompting on long-context clinical summarization and prediction. We examine their ability to synthesize structured and unstructured Electronic Health Records (EHR) data while reasoning over temporal coherence, by re-engineering existing tasks, including discharge summarization and diagnosis prediction from two publicly available EHR datasets. Our results indicate that long context windows improve input integration but do not consistently enhance clinical reasoning, and LLMs are still struggling with temporal progression and rare disease prediction. While RAG shows improvements in hallucination in some cases, it does not fully address these limitations. Our work fills the gap in long clinical text summarization, establishing a foundation for evaluating LLMs with multi-modal data and temporal reasoning.
[…]
tcb@breakable
Large Language Models with Temporal Reasoning for Longitudinal Clinical Summarization and Prediction
@@ 1 Introduction
Electronic Health Records (EHRs) encapsulate a wide range of multi-modal data, such as vital signs, laboratory results, radiology findings, and free text clinical notes Mohsen et al. (2022); Li et al. (2022); Belden et al. (2017). Organized across various timestamps, they reflect the dynamic nature of patient care. In clinical settings, particularly for older patients in intensive care units (ICUs) with multiple encounters, chronic conditions, and complex treatment plans, EHRs become especially lengthy and intricate. Clinicians reviewing these long patient histories face significant cognitive burden, often leading to inaccurate decisions such as diagnostic errors Dymek et al. (2021); Singh et al. (2017). Automated summarization can help improve care continuity and decision-making Dymek et al. (2021); Adams et al. (2021); Gao et al. (2023a); Laxmisan et al. (2012); Pivovarov and Elhadad (2015); Liang et al. (2019), but the volume and complexity of EHRs pose challenges, requiring models that handle diverse and evolving clinical data. As shown in Figure 1, summarizing longitudinal patient trajectories requires extracting temporally ordered events from multi-modal data, with temporal reasoning critical for preserving clinical causality.
@@
Our work evaluates large language models (LLMs) for summarizing complex, longitudinal EHRs, capturing full patient trajectories rather than isolated clinical snapshots. While LLMs have shown promise in clinical NLP tasks Silcox et al. (2024); Wachter and Brynjolfsson (2024); Garcia et al. (2024); Adams et al. (2024); Gao et al. (2022), existing evaluations primarily focus on short-context settings and task-specific fine-tuning. Clinical summarization benchmarks, such as discharge summarization (“Discharge Me”Xu et al. (2024a)) and diagnosis generation from progress notes (“ProbSum”Gao et al. (2023a)), evaluate LLMs on isolated segments of patient records rather than full hospital trajectories. While Retrieval-Augmented Generation (RAG) has been applied to clinical tasks like diagnosis prediction and discharge summarization Xu et al. ; Gao et al. (2023b); Lewis et al. (2020); Myers et al. (2024); Lyu et al. (2024), its effectiveness for long, temporally rich clinical narratives remains unclear.
This work addresses two key gaps: (1) Current evaluations rarely test long-context LLMs in zero-shot settings, leaving their inherent capabilities and limitations in complex medical reasoning unknown; and (2) most clinical benchmarks focus on single time-point summaries, overlooking the demands of longitudinal patient care. We systematically evaluate LLMs on full patient trajectories to assess their ability to process temporally evolving, multi-modal clinical data.
@@
In this paper, we focus on two public EHR datasets: Medical Information Mart for Intensive Care (MIMIC-III) Johnson et al. (2020) and EHRShot Wornow et al. (2023). We evaluate five state-of-the-art LLMs and their RAG variants: Mistral-7B-Instruct-v0.1  Jiang et al. (2023), Llama3-8B-Instruct  AI@Meta (2024), Qwen2.5-7B  Yang et al. (2024), DeepSeek-R1-Distill-Qwen-32B  DeepSeek-AI et al. (2025) and Llama2-13B-chat-hf  Touvron et al. (2023). Our work advances clinical LLM summarization through the following contributions:
@@
  • We reformulate discharge summarization, progress note summarization, and diagnosis classification into new long-context tasks requiring temporal reasoning, aligned with real clinical workflows.
  • Our tasks integrate structured and unstructured data across multiple timestamps, enabling analysis of modality and temporal context effects.
  • We compare direct generation, retrieval-augmented generation and chain-of-thought (CoT) prompting approaches for handling long clinical documents.
In the absence of benchmarks combining multi-modal inputs, temporal structure, and clinical workflow alignment, our study provides a first step toward evaluating LLMs in real-world longitudinal summarization. The results highlight key limitations and suggest directions for more temporally grounded and clinically usable models.
@@ 2 Related Work
Directly processing the long document input can lead to LLMs’ “lost-in-the-middle" problem Liu et al. (2024). As a result, RAG has been the default when handling long clinical documents, evidenced by its superior performance in diagnosis prediction, discharge summarization and information extraction Myers et al. (2024); Lyu et al. (2024); Lopez et al. (2025). Due to the task setup, these works have only investigated single modality EHRs with long input length.
Our work addresses this problem by extending RAG and CoT-based approaches to multi-modal EHR summarization and clinical prediction, incorporating both structured (tabular) and unstructured (clinical notes) data to enhance the completeness and accuracy of patient trajectory summaries. We evaluate how well LLMs and their RAG set ups can capture and preserve temporal and causal relationships across patient’s hospital stay, bridging the gap between isolated document processing and comprehensive longitudinal understanding.
@@ 3.1 MIMIC-III
MIMIC-III is a publicly available dataset comprising de-identified health records from over 40,000 ICU patients. For this study, we focus on a subset of patients with hospital stays exceeding 72 hours to ensure sufficient context for evaluating long-document summarization and temporal reasoning. Unlike prior MIMIC-III studies, our selected cohort emphasizes complex, multi-day ICU stays where accurate summaries are most impactful. This setup allows us to assess LLMs’ ability to handle prolonged and evolving clinical narratives.
We use both structured and unstructured data. Specifically, we extract chart events (vital signs, ventilator settings), lab events (e.g., white blood cell counts), input events (e.g., IV infusions, feedings), and medications. These features reflect key aspects of clinical decision-making and support rich, temporally grounded summarization tasks.
Discharge Summarization
@@
Assessment and Plan (A&P) Generation
@@
The setup of the task is illustrated in Figure 2, which explains the three input methods: the no prior context (Baseline) method uses only the current day’s structured EHR data, the Single-Day Context method adds the previous day’s progress note for limited historical context, and the Multi-Day Context method incorporates all prior progress notes for a more comprehensive patient history. For day $i$, the input includes progress notes up to i1i-1 (depending on the method), but the note for $i$ is excluded and compared to the generated version. This setup reflects the real clinical workflow, where physicians reference prior notes when writing new ones.
@@
Figure 3 shows input lengths across tasks and settings. For discharge summarization, inputs average 500 tokens (24-hour window) and 1,940 tokens (48-hour window). In contrast, the Assessment and Plan (A&P) generation task involves much longer inputs due to the inclusion of prior progress notes and structured data. The No Context baseline starts with a relatively high token count (3,125 tokens on Day 1) but remains the shortest overall. Single-Day Context, which adds the previous day’s note, increases input length to 5,375 tokens. Multi-Day Context, which accumulates all prior notes, shows the steepest token growth—reaching nearly 10,000 tokens by Day 4 and averaging 6,875 tokens.
@@ 3.2 Diagnosis Prediction from EHRShot
EHRShot Wornow et al. (2023) is a longitudinal dataset comprising fully structured EHR data from 6,739 patients at Stanford Medicine, with over 41 million clinical events. In contrast to MIMIC-III, which contains a mix of structured and unstructured data, EHRShot offers a clean, structured-only setting, allowing for controlled evaluation of temporal reasoning in longitudinal patient records.
We include EHRShot to complement the MIMIC-based summarization tasks in two important ways. First, EHRShot focuses on future prediction—each task requires predicting whether a patient will receive one of six diagnoses (e.g., Hypertension, Pancreatic Cancer, Acute MI) within one year after discharge. This differs from our MIMIC tasks, which emphasize retrospective summarization of observed hospital stays. Second, its purely structured nature allows us to isolate models’ abilities to handle temporal patterns in tabular data without the added complexity of clinical language. Together, these contrasts offer a broader assessment of LLM capabilities in modeling temporal EHR information across different modalities and predictive targets.
The six diagnosis tasks vary in prevalence from common conditions like Hyperlipidemia (31.85%) to rarer ones like Celiac (3.46%). Inputs have a mean token length of 1,989 ($\sigma$: 1,270), with a median of 1,851, and an average of 74 unique measurements per patient. More statistics are provided in Appendix A.2.
@@ 3.3 Representing multi-modal, longitudinal patient data
We convert structured EHR data into natural language to enable LLM-based summarization, following prior table-to-text approaches Gao et al. (2024); Yu et al. (2023). As in Table 1, each measurement is verbalized using simple templates (e.g., [Measurement] is [Value][Unit]), grouped by timestamp, and temporally ordered using relative time references. Medications and input events follow a similar format (e.g., “is administered”).
To reduce redundancy from repeated or copy-pasted entries, we deduplicate records across modalities and retain only the most recent entry when multiple identical values appear within a short time window. Clinical notes are filtered similarly. For EHRShot, structured inputs are divided into six hour chunks and repetitive phrases are removed. Appendix A.1.
@@ 4 Methods
We evaluate five LLMs with varying context capacities. LLaMA2-13B and Mistral-7B support up to 4K tokens, while LLaMA3-8B allows for 8K. To test long-context capabilities, we include Qwen2.5-7B and DeepSeek-R1 (32B), both of which support input lengths up to 128K tokens. While we did explore biomedical LLMs (PMC-LLaMa, BioGPT, BioMistral), their performance was extremely poor, so we did not include them in this paper.
@@ 4.1 Approaches
We compare a baseline direct generation, structured RAG and CoT event extraction approach. The key difference between these approaches lies in how they handle long temporal contexts. Direct generation processes the entire input at once but may suffer from the lost-in-the-middle problem, where relevant information buried in long documents is overlooked. In contrast, RAG segments temporal information into retrievable chunks, which helps manage long sequences but may disrupt temporal dependencies between events. For CoT event extraction, the long context is first compressed into a series of clinical events, over which the model then summarizes.
1). Direct Generation.
@@
3). Event Extraction via Chain-of-Thought (CoT).
As an alternative to retrieval or direct generation, we introduce a chain-of-thought (CoT) Wei et al. (2022) prompting approach that incorporates a clinical reasoning step before summarization. LLMs are prompted to first identify temporally ordered clinical events from the input data, which are appended to the original input as structured context for generation. This intermediate step emphasizes salient clinical signals and reduces input noise, helping the model focus on meaningful patterns.
We develop our prompt with guidance from recent studies showing that LLMs perform better on clinical tasks when given clear, structured instructions tailored to the specific goal. For example, Wang et al. (2023) finds that prompts targeting specific event types like symptoms, lab results, treatments, and medical decisions help models focus on clinically meaningful information and improve event detection. Yuan et al. (2023) also shows that when complex tasks are broken into smaller reasoning steps, especially for understanding how events are ordered over time, models tend to produce more consistent and accurate outputs. Figure 4 illustrates the CoT prompt design for A&P generation task.
Chain-of-Thought Prompt for ICU Daily Event Extraction ICU DAILY EVENT EXTRACTION TASK Analyze this structured ICU data by identifying critical clinical events, paying special attention to numerical values and their progression. Only report values showing meaningful change or clinical significance. For repeated values, mention only those demonstrating changes. {Input Text} Identify (with direct references to data points when possible): 1. Major symptoms or changes (new, worsening, improving) – Specify relevant numerical changes. 2. Critical test results (labs, imaging, etc.) – Highlight significant abnormal or normal values. 3. Important treatments or interventions – Clearly link to the preceding clinical data. 4. Significant care team decisions – Support with relevant clinical data. 5. Major medical decisions or diagnoses – Reference pertinent clinical observations. Response Format: ### Day X Key Events ### - [Time] | [Event Description] (Explanation for identifying this event)
Example Model Output ### Day 3 Key Events ### - 2178-02-11 09:50:00 | Blood pressure readings of 88/46 mmHg (This low blood pressure reading is concerning and may require additional fluid resuscitation or adjustment of vasoactive medications.)
Figure 4: Chain-of-Thought prompting template and example output used for ICU daily event extraction from structured EHR data for A&P generation.
We experiment with seven temporal input configurations for event extraction: using only the current day’s data, combining it with previous days (Forward), or with subsequent days (Backward), each spanning 1, 3, or all available days. We apply these settings to a development batch of 20 patients and found that, within each respective group of experiments, the Forward setting using all previous days and the Backward setting using one following day produced the best results. Based on these findings, we report performance under three representative temporal contexts: +0 (Current Day Only, baseline), +n (Forward Context with all prior days), and –1 (Backward Context with one following day).
Experiment settings.
For all LLMs, we run their 8-bit quantized version. We set the output token length as 1,000, but almost all tasks output is significantly shorter than this limit. For RAG setup, we used Langchain Chase (2022) with Faiss Douze et al. (2024) for semantic retrieval in a vector database. We perform hyperparameter tuning with chunk size between [250, 750], top k between [10, 50], chunk overlap between [50,200]. The detailed results of hyperparameter searching are in Appendix A.4. All experiments are run on 2 H100 94GB GPUs.